OpenStudy (anonymous):

the sum of two numbers is 15 and their difference is 5 write a system. let x be the larger number y be the smaller. whats the large number

OpenStudy (anonymous):

sum=x+y diff=x-y

OpenStudy (anonymous):

x+y= 15 x-y=5

OpenStudy (anonymous):

cool what i have on paper. thank u

OpenStudy (anonymous):

so thing is not a answer like im thinking but a equation to get there. right